If you can read the bottom row of your doctor's eye chart, your eye has a resolving power of one arcminute, equal to 1.67E-2 degrees. If this resolving power is diffraction-limited, to what effective diameter of your eye's optical system does this corresponding? Use Rayleigh's criterion and assume that the wavelenght of the light is 555nm.
Answer #1
Concepts and reason

The concept used to solve this question is Rayleigh criterion.

First, convert the value of θ\theta from degree to radians. Use the value of wavelength and θ\theta in the expression of effective diameter.

Fundamentals

The Rayleigh criteria states that the condition of minimum resolution.

The expression of the Rayleigh criteria is given as follows:

1.22λ=Dθ1.22\lambda = D\theta

Here, λ\lambda is the wavelength of light, D is the effective diameter of eye’s optical system, and θ\theta is the small angle in radians.

Convert the value of θ\theta from degree to radians.

θ=0.0167(3.14radians180)=0.000291radians\begin{array}{c}\\\theta = 0.0167^\circ \left( {\frac{{3.14{\rm{ radians}}}}{{180^\circ }}} \right)\\\\ = 0.000291{\rm{ radians}}\\\end{array}

The expression for the Rayleigh criteria is given by,

1.22λ=Dθ1.22\lambda = D\theta

Rearrange the above expression for D as follows:

D=1.22λθD = \frac{{1.22\lambda }}{\theta }

Convert the value of λ\lambda from nm to m.

λ=555nm(109m1.0nm)=555×109m\begin{array}{c}\\\lambda = 555{\rm{ nm}}\left( {\frac{{{{10}^{ - 9}}{\rm{ m}}}}{{1.0{\rm{ nm}}}}} \right)\\\\ = 555 \times {10^{ - 9}}{\rm{ m}}\\\end{array}

Substitute 555×109m555 \times {10^{ - 9}}{\rm{ m}}for λ\lambda and 0.000291 radians for θ\theta in the equationD=1.22λθD = \frac{{1.22\lambda }}{\theta }.

D=1.22(555×109m)0.000291radians=2.33×103m\begin{array}{c}\\D = \frac{{1.22\left( {555 \times {{10}^{ - 9}}{\rm{ m}}} \right)}}{{0.000291{\rm{ radians}}}}\\\\ = 2.33 \times {10^{ - 3}}{\rm{ m}}\\\end{array}

Ans:

The value of the effective diameter is 2.33×103m2.33 \times {10^{ - 3}}{\rm{ m}}.